television for $9000 down and monthly payments of $500. When he visited the lot, he found it was not something he wanted to own. After 40 months he was finally able to sell the lot. The new owner assumed the balance of the loan on the lot and paid Bill $19,000. What rate of return did Bill receive on his investment? e project, which had a projected life of 5 years, was terminated early. Compute the interest rate. 7A-12 The Year Cash Flow 0 -$5000 1500 1500 2 7A-13 Compute the rate of return on the investment. Year Cash Flow $5000 7500 1500 -6000 0 7A-14 Consider the following cash flow. Year Cash Flow 1 00 Please only do 7A-13 use 6% for financing rate and 12% for the investing rate
Answer #1

Since, there are changes in the cash flows from positive to negative in the given question, we will have to calculate the Modified Internal Rate of Return on Investment with the use of following formula:

Modified Internal Rate of Return on Investment = (Sum of Future Value of Cash Inflows at Reinvestment Rate/-Present Value of Cash Outflows at Finance Rate)^(1/Years) - 1

Using the values provided in the question, we get,

Sum of Future Value of Cash Inflows at Reinvestment Rate = Cash Flow Year 1*(1+Reinvestment Rate)^2 + Cash Flow Year 2*(1+Reinvestment Rate)^1 = 7,500*(1+12%)^2 + 1,500*(1+12%) = $11,088

Present Value of Cash Outlows at Finance Rate = -5,000 - 6,000/(1+6%)^3 = -$10,037.71

Substituting these values in the above formula for MIRR, we get,

Modified Internal Rate of Return on Investment = (11,088/-(-10,037.71))^(1/3) - 1 = 3.37% (answer)

The rate of return on investment would be 3.37% in the given case.